Five friends drove at an average rate of 50 miles per hour to a weekend retreat. On the way home, they took the same route but averaged 65 miles per hour. What was the distance between home and the retreat if the round trip took 10 hours?
Answer by richwmiller(17219)

50*x=65(10-x)
x = 130/23
50*130/23=d
d=282 14/23=282.6 miles